On september 11, 2003, the Minneapolis Star Tribune compared the price of Glad wrap to Press 'n Seal. A roll of Glad wrap sells for $2.99, A roll of Press n' Seal sells for $3.79 or 5 cents a squeare foot. Press's Seal for 10/3 times as much per square foot as Glad wrap. What is the price of Glad wrap per square foot?
Answer by stanbon(48568)

A roll of Glad wrap sells for $2.99,
A roll of Press n' Seal sells for $3.79 or 5 cents a square foot.
Press's Seal for 10/3 times as much per square foot as Glad wrap.
What is the price of Glad wrap per square foot?
--------------
EQUATION:
PS =(10/3)GW
5 cents/sq. ft = (10/3)GW
GW = 5/(10/3) = = 15/10 = 1.5 cents/sq.ft
===============
Cheers,
Stan H.